OpenStudy (luigi0210):

Sequences:

OpenStudy (luigi0210):

Converges or Diverges? \[\LARGE \frac{(-1)^n}{2\sqrt{n}}\]

OpenStudy (anonymous):

alternating sequences converge if... 1. Its terms decrease in magnitude 2. The terms converge to 0.

OpenStudy (anonymous):

have you tried the ratio test?

hartnn (hartnn):

Alternating series test[edit] This is also known as the Leibniz criterion. If \(\sum_{n=1}^\infty a_n\) is a series whose terms alternative from positive to negative, and if the limit as n approaches infinity of a_n is zero and the absolute value of each term is less than the absolute value of the previous term, then \(\sum_{n=1}^\infty a_n \) is convergent. source : wikipedia

OpenStudy (nincompoop):

try solving

OpenStudy (luigi0210):

\[\LARGE \lim_{n \rightarrow \infty}~\frac{(-1)^n}{2\sqrt{n}}\] As a limit?

OpenStudy (nincompoop):

how else laughing out loud

OpenStudy (luigi0210):

You'd be surprised >_>

OpenStudy (luigi0210):

Wouldn't that make it an indeterminate at the top and infinity at the bottom?

OpenStudy (nincompoop):

try what hartnn suggested

OpenStudy (anonymous):

i think its conv

OpenStudy (sidsiddhartha):

do know leibnitz theorem for a alternating series?

OpenStudy (anonymous):

it is converge but heres what i do

OpenStudy (luigi0210):

Yup, each terms gets smaller and smaller.. \(\Large |-1/2|, |1/2\sqrt{2}|, |-1/2\sqrt{3}|, |1/4|,... \)

OpenStudy (luigi0210):

I'm technically not suppose to know any "theorems" by this chapter..

OpenStudy (nincompoop):

you don't

OpenStudy (luigi0210):

Or section

OpenStudy (nincompoop):

just use whatever else you know

OpenStudy (anonymous):

when n is odd lim =0 when its even lim =0

OpenStudy (sidsiddhartha):

a series of the form \[un=\sum_{n=1}^{\infty}(-1)^{n}un\] converges if 1){un} is monotonic decreasing and 2)\[\lim_{n \rightarrow \infty}un=0\]

OpenStudy (sidsiddhartha):

check if the conditions satisfies

OpenStudy (anonymous):

its not monotonic.... u can devide it into seq you would have one decrease and one increase

OpenStudy (sidsiddhartha):

it is convergent the conditions are satisfying

OpenStudy (anonymous):

just devide it into two subsequences \(\LARGE \lim_{n \rightarrow \infty}~\frac{ -1 }{2\sqrt{2n+1}}\) \(\LARGE \lim_{n \rightarrow \infty}~\frac{1 }{2\sqrt{2n}}\) if lim 1=lim 2 = C s.t C\(\neq 0\) then its conv.

OpenStudy (anonymous):

i made a typo sry c\(\neq \infty \)

OpenStudy (sidsiddhartha):

\[u_n-u_{n-1}>0\] and also \[\lim_{n \rightarrow \infty}\frac{ (-1)^n }{ 2\sqrt(n)}=0\]

OpenStudy (anonymous):